# Towards Geo-Culturally Grounded LLM Generations

### v2

@@ Methods.
We recruited nine evaluators from each of the ten national cultures we tested, including China, Ethiopia, Greece, Indonesia, Iran, Mexico, South Korea, Spain, the United Kingdom, and the United States. Each evaluator was shown a prompt relevant to their national culture, along with responses to that prompt from each of the three strategies evaluated: vanilla, selective KB-grounding, and search-grounding. For each strategy, a response was randomly selected from the three responses generated. Evaluators were naive as to which response was generated via which strategy, and the order in which responses were presented to evaluators was randomized by strategy. Evaluators were then asked to rate with justification, on a scale from 0 to 4, how culturally informed each response was. Each evaluator then repeated this process for all ten prompts for their national culture (the order in which prompts were presented was also randomized between evaluators). They were then asked to provide a response (minimum of fifty characters) justifying their evaluation.
@@ A.8 Sample Question and Responses from the Human Evaluation Experiment
Sample generated texts in Table 8 are the responses from the Gemini model for the question: "Tell me a story in Mexico in which a group of people of varying ages eat together and all guests behave in a socially acceptable way."
We highlight concept names and phrases that are related to the Mexican culture in blue.

### v3

@@ Methods.
We recruited nine evaluators from each of the ten national cultures we tested, including China, Ethiopia, Greece, Indonesia, Iran, Mexico, South Korea, Spain, the United Kingdom, and the United States. Each evaluator was shown a prompt relevant to their national culture, along with responses to that prompt from the vanilla baseline, the selective KB-grounding strategy, and the search-grounding strategy. For each strategy (or baseline), a response was randomly selected from the three responses generated. Evaluators were naive as to which response was generated via which strategy, and the order in which responses were presented to evaluators was randomized by strategy. Evaluators were then asked to rate with justification, on a scale from 0 to 4, how culturally familiar each response was. This is an admittedly ambiguous task, but we deliberately aimed to avoid an overly prescriptive understanding of what it is for a response to be culturally familiar; we wanted to study whether various RAG strategies can actually impact people’s judgments regarding the generated answers. Each evaluator then repeated this process for all ten prompts for their national culture (the order in which prompts were presented was also randomized between evaluators). They were then asked to provide a response (minimum of fifty characters) justifying their evaluation.
We deliberately structured our study so as to avoid a need for inter-evaluator calibration. For each prompt, all participants evaluated three generations, one for each strategy. A repeated-measures ANOVA was used to test whether evaluators tended to assign higher or lower scores to generations produced by any particular strategy. Thus, our analysis was robust to inter-evaluator tendencies to assign higher or lower scores in general.
@@ A.8 Sample Question and Responses from the Human Evaluation Experiment
Sample generated texts in Table 8 are the responses from Gemini for the question: "Tell me a story in Mexico in which a group of people of varying ages eat together and all guests behave in a socially acceptable way." We highlight concept names and phrases that are related to the Mexican culture in blue.
@@
A.9 Statistics of the Retrieved Documents
Fig. 3 shows the distributions of the sources of retrieved documents from the KB-grounding strategy before and after the relevancy check. Fig. 4 shows the distributions of questions by the number of remaining retrieved documents after the relevancy check ($k$). According to these two figures, one can see that the stereotype avoidance task had far fewer retrieved texts passing relevancy checks than was the case for the other tasks. As shown in Fig. 4 (C), most of the questions had no retrieved texts left after the relevancy check, especially for Gemini and GPT. This is a good result, as it demonstrates that the models typically treat information as irrelevant for predicting the nationality given only the stereotypical attribute. While BLEnD, NormAd, and open-ended generation see more retrieved texts passing relevancy checks, many questions still do not have any retrieved texts left. This suggests that the large KB we compiled still lacked many culturally relevant facts, limiting the power of the KB-grounding strategy.
Ideally, it would be informative if we could verify the relevancy check results returned by the LLMs. However, we anticipate two major challenges. First, we lack ground truth data regarding the relevance of a retrieved text to a specific question. Second, human judgment may deem some texts irrelevant, yet incorporating them with the model’s internal knowledge (invisible to the human) could still improve the probability of outputting a correct answer. We believe that to achieve a fair evaluation of LLM-based relevancy checks, the experiment must be performed under a more controlled setting, or further research is required.
Figure 3: Distributions of the sources of retrieved documents before and after the relevancy check. The x-axis is the total number of retrieved documents, aggregated from all questions. The y-axis is the LLM that performed the relevancy check where ‘None’ shows the source distribution before the relevancy check was applied.
Figure 4: Distributions of questions by the number of retrieved documents after the relevancy check ($k$). Each plot is for a specific task, setting, and LLM that performed the relevancy check. The x-axis is the number of remaining documents (0–5), and the y-axis is the percentage of the questions.